how do you solve :
2x-6y=-16
5x+7y=-18

Answer by checkley71(8403)

multiply the first equation by 5 10x-30y=-80
now multiply the second equation by -2 -10x-14y=36
now add these two equations
10x-30y=-80
-10x-14y=36
--------------------------
-44y=-44
y=-44/-44
y=1 answer
2x-6*1=-16
2x-6=-16
2x=-16+6
2x=-10
x=-10/2
x=-5 answer.
proof
5*-5+7*1=-18
-25+7=-18
-18=-18